OpenStudy (anonymous):

log10[x^n/100y] simplify

OpenStudy (anonymous):

Use the rules: \[\begin{align} \log(ab)&=\log(a)+\log(b)\\ \log(a^b)&=\log(a)\cdot\log(b)\\ \log\left(\frac{a}{b}\right)&=\log(a)-\log(b) \end{align}\]

OpenStudy (anonymous):

\[\log(x^n)-\log(100y)\] is a start then \(n\log(x)-\log(100)-\log(y)\) and finish by finding \(\log(100)\)

OpenStudy (anonymous):

thank you!!!!